TheKnightKing8273:

The area of a regular polygon is 35cm. What is the area of a regular octagon with sides eight times as long? A-280cm B-1600cm C-2440cm D-350cm

TheKnightKing8273:

Please help!!!!

jhonyy9:

ATTENTION !!! area not can be cm just cm squared so cm^2

supie:

All we have to do is use the area of an octagon formula to find the length of the sides but we only need to find the length of one side.

supie:

So, what we would have to do is \[A = 2 ( 1 + √2) s^2\]\[35 = 2 ( 1 + √2) s^2\]\

supie:

Now all we have to do is use the area of an octagon formula to find the area of the 2nd octagon by ()ing the side length above by five the formula \[2a^2(1+√2)\]

supie:

So, all we have to do is \[A = 2 ( 1 + √2) s^2\]\[A = 2 ( 1 + √2) (2.57 • 5)^2\]\[A = 736.44 cm^2\]